OpenStudy (anonymous):

i need to solve 6x-10x+5=-15

OpenStudy (anonymous):

multiple choice?

OpenStudy (anonymous):

rearrange the equation so that all x-terms with their coefficients are on one side and the other numbers are on the other. you'll get 6x-10x = -15 -5 then, simplify each side: (6-10)x = -15-5 .... -4x = -20 then divide both sides by -4... x = -20/-4 so, x =5